OpenStudy (anonymous):

The LCM of 21w and 9w^2 is 63w^2, correct?

OpenStudy (deadpoolroxz):

21*9?

OpenStudy (anonymous):

No..

OpenStudy (deadpoolroxz):

im sorry i do not know wesh i could help peace im out

OpenStudy (jdoe0001):

well, what factors does 21 have? what about 9? what factors does it have? do a prime factoring of them, let's see what factors they have then

OpenStudy (anonymous):

@jdoe0001 21: 3*7 9: 3*3 3*3*7=63

OpenStudy (jdoe0001):

yeap the prime factoring shows, that 3 * 3 * 7 are it :)

OpenStudy (jdoe0001):

\(21={\color{brown}{ 3}}\cdot {\color{brown}{ 7}} \\ \quad \\ 9={\color{brown}{ 3 }}\cdot 3 \\ \quad \\ LCM = {\color{brown}{ 3\cdot 3\cdot 7}}\) since, that will include, 21, as well as 9 IN IT

OpenStudy (anonymous):

Okay, but what about the variables?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

ohhh hmmm well, you have \(w\ and\ w^2\) it has to be \(w^2\) because that includes "w" in it :)

OpenStudy (deadpoolroxz):

yea ik i havent i wasnt payin attention i was doing homework

OpenStudy (anonymous):

Lol, nice.

OpenStudy (deadpoolroxz):

dude fan me